OpenStudy (anonymous):

Write with expression with base 2 : 4^6

OpenStudy (mathmate):

4^6 =(2^2)^6 =2^12 by the laws of exponents (a^b)^c = a^(bc)

OpenStudy (anonymous):

@mathmate how would i write 14 with a base of 2

OpenStudy (mathmate):

14 cannot be written with a base of 2 (only). Can you post the complete question?

OpenStudy (anonymous):

that is the complete question

OpenStudy (mathmate):

2^3+2^2+2^1 or 2^4-2^1